A jaw crusher, with a closed setting of 3 in., produces 200 tph of crushed stone. Determine the number f 


1
Expert's answer
2021-10-05T17:53:55-0400

Complete questions:A jaw crusher, with a closed setting of 3 in., produces 200 tph of crushed stone. Determine the number of tons per hour produced in each of the following ranges.

"In\space excess\space 2\frac{1}{2}in\\
between\space 2\frac{1}{2}in\space and \space 1\frac{1}{2}in\\
Between\space 1\frac{1}{2}in\space \frac{1}{4}in\\
Less\space than\space \frac{1}{4}in"

The total output 150 tph of crushed stone.

Solution

We obtain percentage in size range and amount output in size range.

We use crusher setting of 4 to obtain percentage for each size.

Formula

Amount produced in size range=% in size range× total output of crusher.


"In\space excess\space 2\frac{1}{2}in"  in the % size range

"100-56=44"

Amount produced "0.44\u00d7150=66"


"In\space between \space 2\frac{1}{2}in \space and \space 1\frac{1}{2}in" ​ in the % size range

"56-36=20"

Amount produced "0.2\u00d7150=30"


"In\space between \space 2\frac{1}{2}in \space and \space 1\frac{1}{2}in" ​ in the % size range

"36-11=25"

Amount produced "0.25\u00d7150=37.5"



"less\space than\space 2\frac{1}{2}in" ​ in the % size rang

"11-0=11"

Amount produced "0.11\u00d7150=16.5"